OpenStudy (anonymous):

There are 30 people in student council that are running for the offices of president and vice president. In how many different ways can those offices be assigned?

OpenStudy (anonymous):

1/30*1/30= your answer

OpenStudy (anonymous):

There are 30 people in student council that are running for the offices of president and vice president. In how many different ways can those offices be assigned? 59 ways 60 ways 870 ways 900 ways

OpenStudy (anonymous):

First, you fill one position with a pool of 30 applicants. So you have 30 ways to fill that position. Next, the first position is taken, so the next position has only 29 applicants. So there are 29 ways to fill that. So the total ways you can fill TWO positions is 30 * 29 = 870 ways. What about 3? Keep going! Now there are 28 students for the third application. See where this is going? This is called a factorial. 30-factorial is written as '30!' and: 30! = 30 * 29 * 28 * ... * 3 * 2 * 1 = a very large number

OpenStudy (anonymous):

if you do it the way i gave you its 900

OpenStudy (anonymous):

trust me i passed my whole unit using that method